I need help, please. I have to solve this problem using an algebraic equation.
Heather has twice as many dimes as nickels and two more quarters than nickels. There were 22 coins in all.
a. How many quarters does she have?
b. The value of the coins is $5.50. How many quarters does she have?

Answer by Paul(988)

2x+x+2+x=22
4x=20
x=5
5+2=7
a) she has 7 quarters


2(10x)+x+25(2+x)=550
20x+x+50+25x=550
46x=500
x=10 (rounded)
10+2=12
B) 12 quarters